OpenStudy (anonymous):

What is the solution to the equation? 4(3 + 7) = 4 • x + 4 • 7 A. x = 3 B. x = 4 C. x = 7 D. x = 10

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

Which one is rught

OpenStudy (anonymous):

right8

OpenStudy (anonymous):

i need 1

OpenStudy (anonymous):

i only can pick 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its x=3

OpenStudy (anonymous):

b/c 3+7=10 4 x 10+40 factor out a 4 from (4*x) and (4*7) then 40/4 =10 10=x+7 10-7=3 x=3

OpenStudy (anonymous):

welll i failed because of you @shaadykhamlichi